Given n(x) = 6x^2 + 11, find n(-3).

Answer by checkley77(12569)

n(x) = 6x^2 + 11, find n(-3).
n(-3)=6*(-3)^2+11
n(-3)=6*9+11
n(-3)=54+11
n(-3)=65 ans.